# LLM $\TITLELLM for Large-Scale Optimization Model Auto-Formulation$: A Lightweight Few-Shot Learning Approach

### v1

@@ 5.4 Performance on Network Planning
Table 12: Execution accuracy of different approaches on Air-NRM-NP
Model Execution accuracy
Gemini 3 Pro 85.71%
GPT-5.2 55.56%
LEAN-LLM-OPT (GPT-4.1) 92.06%
LEAN-LLM-OPT (gpt-oss-20B) 82.54%
LEAN-LLM-OPT (GPT-4.1, tools-only) 30.16%
LEAN-LLM-OPT (GPT-4.1, workflow-only) 0.00%
LEAN-LLM-OPT (GPT-4.1, base model-only) 0.00%
We next turn to Air-NRM-NP and examine the execution accuracy and revenue loss incurred when deploying LEAN-LLM-OPT to plan the airline network. In Air-NRM-NP, each instance requires formulating and solving an optimization model that jointly determines the number of seats allocated to each fare class and selects a subset of flights to maximize revenue. The execution accuracies of different methods are reported in Table 12. We observe that LEAN-LLM-OPT (GPT-4.1) continues to achieve the highest execution accuracy, with Gemini 3 Pro performing second best. Moreover, LEAN-LLM-OPT (gpt-oss-20B) attains execution accuracy comparable to Gemini 3 Pro. Given that gpt-oss-20B has substantially fewer parameters than the other models considered, these results suggest that LEAN-LLM-OPT can help elicit and steer the underlying model’s reasoning in this setting.
@@
Figure 7 reports the optimality gaps of different methods on Air-NRM-NP as the problem size varies. As shown, the optimality gap of LEAN-LLM-OPT (GPT-4.1) remains consistently below 14%, whereas the tools-only counterpart typically yields gaps exceeding 55% for most instances. The optimality gap of LEAN-LLM-OPT (gpt-oss-20B) fluctuates but stays below 24% in all but two cases. This trend is similar to Gemini 3 Pro, whose optimality gap remains below 17.00% except for two cases. In contrast, LEAN-LLM-OPT (GPT-4.1, tools-only) and GPT-5.2 often exhibit optimality gaps above 40%, and in some cases exceed 80%.
@@ 6 Concluding Remarks
  • First, our results indicate that our interpretable design can substantially enhance the modeling ability of earlier-generation LLMs (e.g., GPT-4.1) and smaller open models (e.g., gpt-oss-20B), enabling them to formulate large-scale optimization models without requiring extensive additional training resources. This suggests that, in specialized domains, incorporating practices into the system design can be a particularly efficient approach;
  • Next, our results suggest that older-generation models and less powerful open-source models may still have considerable potential for solving challenging problems, provided that their reasoning is appropriately elicited and guided. Given that such models often offer lower cost and/or privacy advantages, developing systematic methods for steering them toward competitive performance warrants further study;
  • Finally, our results suggest that, for complex tasks, structured and step-by-step workflows can help LLM agents achieve more competitive performance. In our work, we introduce structure primarily through semantic instructions. Given that workflow and process design has long been a central topic in operations research and operations management (Cachon and Terwiesch 2024), our results indicated that core knowledge in operations has the potential to guide better design of LLM-based systems. It is thus interesting to explore how to design more verifiable mechanisms to further improve performance and reliability systematically. One promising direction is to follow the assured autonomy framework proposed in Dai et al. (2025), which ensures formal structure, constraints satisfaction, accountability, and tail-risk control through an operations research lens.
@@ A.1 Application Domains of the Problem Instances
  • Facility Location Problem (FLP): If the dataset includes a list of potential facility locations with associated fixed opening costs, customer demand quantities, and transportation costs from facilities to customers, this problem can be classified as FLP.
@@ Appendix B Problem Classification Standard
  • NRM (Network Revenue Management): Managing resources across a network to maximize revenue (Talluri and Van Ryzin 2006).
  • RA (Resource Allocation): Distributing and managing limited resources among competing activities to optimize an objective (Winston 2004).
  • TP (Transportation Problem): Minimizing the cost of shipping goods from a set of sources to destinations (Winston 2004).
  • FLP (Facility Location Problem): Determining the optimal placement of facilities to minimize costs while meeting demands from clients(Daskin 1997).
  • AP (Assignment Problem): Optimally assigning agents to tasks on a one-to-one basis (Winston 2004).
@@ Appendix F Type-Tailored Retrieved Data Format
In this section, we present several examples for the retrieved data for one assignment problem (AP), one transportation problem (TP), and one facility location problem (FLP) to complement the type-tailored workflow in Section 3.2.1. The first is an AP where the data can be most efficiently stored as a matrix.
@@
The next example is the data format for transportation problem (TP) where the demand an supply data are stored in a table and the cost data between demand and supply locations is stored in a matrix.
@@ 61 HumanMessage(content=prompt)
Appendix H Example Problem Instance in Air-NRM-CA

### v2

@@ 5.4 Performance on Network Planning
Table 13: Modeling accuracy of different approaches on Air-NRM-NP
Model Modeling accuracy
Gemini 3 Pro 85.7%
GPT-5.2 55.6%
LEAN-LLM-OPT (GPT-4.1) 92.1%
LEAN-LLM-OPT (gpt-oss-20B) 82.5%
LEAN-LLM-OPT (GPT-4.1, tools-only) 30.2%
LEAN-LLM-OPT (GPT-4.1, workflow-only) 0%
LEAN-LLM-OPT (GPT-4.1, base model-only) 0%
We next turn to Air-NRM-NP and examine the modeling accuracy and revenue loss incurred when deploying LEAN-LLM-OPT to plan the airline network. In Air-NRM-NP, each instance requires formulating and solving an optimization model that jointly determines the number of seats allocated to each fare class and selects a subset of flights to maximize revenue. The modeling accuracies of different methods are reported in Table 13. We observe that LEAN-LLM-OPT (GPT-4.1) continues to achieve the highest modeling accuracy, with Gemini 3 Pro performing second best. Moreover, LEAN-LLM-OPT (gpt-oss-20B) attains modeling accuracy comparable to Gemini 3 Pro. Given that gpt-oss-20B has substantially fewer parameters than the other models considered, these results suggest that LEAN-LLM-OPT can help elicit and steer the underlying model’s reasoning in this setting.
@@
Figure 7 reports the optimality gaps of different methods on Air-NRM-NP as the problem size varies. As shown, the optimality gap of LEAN-LLM-OPT (GPT-4.1) remains consistently below 14%, whereas the tools-only counterpart typically yields gaps exceeding 55% for most instances. The optimality gap of LEAN-LLM-OPT (gpt-oss-20B) fluctuates but stays below 24% in all but two cases. This trend is similar to Gemini 3 Pro, whose optimality gap remains below 17% except for two cases. In contrast, GPT-5.2 often exhibit optimality gaps above 40%, and in some cases exceed 80%.
@@ 6 Concluding Remarks
  • Generalizability of the Agentic Workflow Approach: Although our main focus is optimization modeling, the idea of agentic workflow may extend beyond it. In many domains, valuable and structured expertise has accumulated over time. While such expertise is rarely one-size-fits-all, LLMs provide a powerful mechanism for making these practices more customizable and flexible. By translating domain expertise into (the skeleton of the) workflows while allowing LLMs to handle inherently unstructured components, we may be able to improve both effectiveness and efficiency across a broad range of applications.
  • Stimulate LLM Capabilities: Next, our results suggest that older-generation models and less powerful open-source models may still have considerable potential for solving challenging problems, provided that their reasoning is appropriately elicited and guided. Given that such models often offer lower cost and/or privacy advantages, developing systematic methods for steering them toward competitive performance warrants further study.
  • Enforceable and Controllable Structure: In our work, we introduce structure and workflow primarily through semantic instructions. However, it has been widely-recognized that LLM’s behavior is inherently probabilistic and can make the system’s behavior unstable. It is thus interesting to explore how to design more enforceable and controllable mechanisms to further improve performance and reliability systematically. One promising direction is to follow the assured autonomy framework proposed in Dai et al. (2025), which ensures formal structure, constraints satisfaction, accountability, and tail-risk control through an operations research lens.
@@ A.1 Application Domains of the Problem Instances
  • Facility Location Problem: If the dataset includes a list of potential facility locations with associated fixed opening costs, customer demand quantities, and transportation costs from facilities to customers, this problem can be classified as FLP.
@@ Appendix B Problem Classification Standard
  • Network Revenue Management: Managing resources across a network to maximize revenue (Talluri and Van Ryzin 2006).
  • Resource Allocation: Distributing and managing limited resources among competing activities to optimize an objective (Winston 2004).
  • Transportation Problem: Minimizing the cost of shipping goods from a set of sources to destinations (Winston 2004).
  • Facility Location Problem: Determining the optimal placement of facilities to minimize costs while meeting demands from clients(Daskin 1997).
  • Assignment Problem: Optimally assigning agents to tasks on a one-to-one basis (Winston 2004).
@@ Appendix F Type-Tailored Retrieved Data Format
In this section, we present several examples for the retrieved data for one AP, one TP, and one FLP to complement the type-tailored workflow in Section 3.2.1. The first is an AP where the data can be most efficiently stored as a matrix.
@@
The next example is the data format for TP where the demand an supply data are stored in a table and the cost data between demand and supply locations is stored in a matrix.
@@ 61 HumanMessage(content=prompt)
Appendix H Example for Structurally Incorrect Model with a Correct Optimal Solution
In this section, we present several representative examples from Large-Scale-OR where different models produce correct optimal solutions despite formulating the underlying optimization problems incorrectly. Table 6 summarizes the modeling accuracy of different models on Large-Scale-OR. The cases collected in this section complement the quantitative results by showing how structurally incorrect models may still return the same optimal solutions as the ground-truth formulations.
NRM (Generated by LEAN (GPT-4.1)) Query: The mobile device retailer offers a portfolio of products with heterogeneous regional demand. Each product $i$ yields revenue $R_{i}$ and has a fixed initial inventory level $I_{i}$, with no restocking allowed during the sales horizon. Customer orders arrive stochastically, and the retailer determines an integer fulfillment quantity $x_{i}$ for each product in order to maximize total expected revenue. Demand constraints impose an upper bound $D_{i}$ on feasible fulfillment. The decision problem is to optimally allocate limited inventory across products under both inventory and demand limits. Ground-Truth Model (Label). Objective: $\max\sum_{i=1}^{71}R_{i}x_{i}$ Constraints: $x_{i}\leq I_{i},\hskip 18.49988pt$ $i=1,\dots,71,$ $x_{i}\leq D_{i},\hskip 18.49988pt$ $i=1,\dots,71,$ $x_{i}\in\mathbb{Z}^{+}.$ Model Generated by LEAN (GPT-4.1). Objective: $\max\sum_{i=1}^{71}R_{i}x_{i}$ Constraints: $x_{i}\leq I_{i},\hskip 18.49988pt$ $i=1,\dots,71,$ $x_{i}\in\mathbb{Z}^{+},\hskip 18.49988pt$ $i=1,\dots,71.$ Key Difference. The generated model omits the demand constraints $x_{i}\leq D_{i}$, resulting in a strictly larger feasible region. However, if the optimal solution $x^{\star}$ of the generated model satisfies $x_{i}^{\star}\leq D_{i}\quad\forall i,$ then $x^{\star}$ also satisfies $x_{i}^{\star}\leq\min\{I_{i},D_{i}\},$ Therefore, although the numerical optimal solution coincides with that of the true formulation for this instance, the agreement is incidental.
Others (Generated by ORLM) Query. You are a nutritionist tasked with designing a one-day meal plan using various available foods to minimize the total cost of the menu while meeting specific nutritional guidelines. Per-serving nutrient and price information is provided in cost.csv, which lists Calories (kcal), Protein (g), Fat (g), Vitamin C (mg), and Cost (USD) for each food. Your plan must supply at least 2000 Calories, at least 50 g of protein, at least 60 mg of vitamin C, and no more than 70 g of fat. Portions may be fractional (e.g., 0.5 serving). Determine the optimal number of servings of each food that satisfies these requirements at the lowest possible cost, and clearly formulate the corresponding optimization model. Ground-Truth Model (Label). Decision variables. The labeled model introduces one continuous decision variable for every food item appearing in the dataset: $x_{i}\geq 0,\quad\forall i\in\mathcal{F}.$ The full set $\mathcal{F}$ consists of multiple categories: Grains: 50 items Dairy: 50 items Meat: 50 items Fruit: 50 items Vegetables: 36 items Staples (e.g., Oatmeal, Milk, Egg, Banana): 4 items yielding a total of $|\mathcal{F}|=240\text{ decision variables.}$ Objective: $\min\sum_{i\in\mathcal{F}}c_{i}x_{i}$ Constraints: $\sum_{i\in\mathcal{F}}a_{i}^{\text{cal}}x_{i}$ $\geq 2000,$ $\sum_{i\in\mathcal{F}}a_{i}^{\text{pro}}x_{i}$ $\geq 50,$ $\sum_{i\in\mathcal{F}}a_{i}^{\text{vitC}}x_{i}$ $\geq 60,$ $\sum_{i\in\mathcal{F}}a_{i}^{\text{fat}}x_{i}$ $\leq 70.$ Model Generated by ORLM. Decision variables. The ORLM-generated formulation introduces decision variables for multiple food categories present in cost.csv. However, for each category, only a restricted subset of items is instantiated. Concretely, the generated model includes: Grains: 25 items (out of 50), Dairy: 25 items (out of 50), Meat: 25 items (out of 50), Fruit: 25 items (out of 50), Vegetables: 25 items (out of 36), Staples (Oatmeal, Milk, Egg, Banana): 4 items. As a result, the effective decision set $\widehat{\mathcal{F}}\subset\mathcal{F}$ satisfies $|\widehat{\mathcal{F}}|\approx 129,$ which is substantially smaller than the full labeled model with $|\mathcal{F}|=240.$ Objective and constraints. The objective function and nutritional constraints are formulated in the same algebraic form as the label model, but are evaluated only over the reduced set $\widehat{\mathcal{F}}$, rather than the complete food universe. Key Difference. The labeled formulation optimizes over the full food universe $\mathcal{F}$, whereas the ORLM model restricts decisions to a much smaller subset $\widehat{\mathcal{F}}\subset\mathcal{F}$. This dimensionality reduction fundamentally alters the feasible region and the optimization problem itself. Although the two models are not structurally equivalent, their optimal solutions may coincide under particular parameter realizations—for instance, when the omitted food items are not selected in any cost-minimizing feasible solution. Such agreement is instance-specific and does not imply correctness of the generated formulation.
NRM (Generated by Gemini 3 Pro) Query. The supermarket offers a variety of products, with associated revenue data provided in the ’Revenue’ column. Each product has an independent demand stream. The retailer aims to maximize total expected revenue by focusing on the initial inventory of all the products. The initial inventory levels for the products are provided in the ’Initial Inventory’ column. During the sales horizon, no restocking is allowed, and there are no in-transit inventories. Throughout the sales horizon, customer inquiries for products arrive sequentially, following a Poisson process, with demand information specified in the ’Demand’ column. The decision variables, represented by $x_{i}$, indicate the number of customer requests the company plans to fulfill for each product $i$, with each $x_{i}$ being a positive integer. Ground-Truth Model (Label). Decision variables. $x_{i}\in\mathbb{Z}_{\geq 0},\quad i=1,\dots,109.$ Objective: $\max\sum_{i=1}^{109}R_{i}x_{i}.$ Constraints. $x_{i}\leq I_{i},\qquad i=1,\dots,109,$ $x_{i}\leq D_{i},\qquad i=1,\dots,109.$ Implied feasible region. $0\leq x_{i}\leq\min\{I_{i},D_{i}\},\quad i=1,\dots,109.$ Model Implied by Gemini 3 Pro (Code). Decision variables. $x_{i}\in\mathbb{Z}_{\geq 0},\quad i\in\mathcal{P},$ where $\mathcal{P}$ is the product list read from the CSV (one variable per row; in this instance it matches 109 products). Objective. Gemini replaces $\sum_{i}R_{i}x_{i}$ with a stochastic expected-sales objective: $\max\sum_{i\in\mathcal{P}}R_{i}\cdot\mathbb{E}\!\left[\min\{x_{i},\ \mathrm{Pois}(\mu_{i})\}\right],$ Constraints. $0\leq x_{i}\leq I_{i},\qquad i\in\mathcal{P},$ $\sum_{i\in\mathcal{P}}x_{i}\leq C,$ with a hard-coded total capacity $C=150000$. Key Differences. Objective mismatch. the label model is linear in $x_{i}$, while the generated code optimizes a Poisson expected-sales function using a piecewise-linear approximation. Missing demand upper bounds. the label includes $(\mathrm{Dem})\ x_{i}\leq D_{i}$ for every product, but the generated code does not enforce these per-product demand limits as constraints. Extra global capacity constraint. the generated code adds $(\mathrm{Cap})\ \sum_{i}x_{i}\leq C$ with $C$ fixed to 150000, which is not present in the label formulation. The two formulations are not structurally equivalent. Nevertheless, their optimal solutions can coincide for certain datasets—for example, when the optimizer under the generated model happens to satisfy $x_{i}^{\star}\leq D_{i}$ for all $i$ and the added capacity constraint is non-binding.
Mixture6 (Generated by LEAN (gpt-oss-20b)) Query: A logistics company needs to schedule three trucks daily to transport goods and meet customer demand across four time periods. Each truck has specific startup and transportation costs. The customer demand for the four time periods is [1,500 kg, 2,000 kg, 1,800 kg, 1,000 kg]. The company must follow these operating rules: trucks have minimum runtime and cooldown requirements. Once a truck is started, it must run continuously for at least 2 hours before being shut down. If turned off, it must remain idle for at least 1 hour before it can be restarted. there are limits on load fluctuations. For each individual truck, the change in transported weight between two adjacent periods cannot exceed 300 kg. Additionally, the company must always maintain a 10% spare capacity buffer for emergency orders. This means the total transported weight in any period must not exceed 90% of the combined maximum capacity of the trucks active in that period. The goal is to find the lowest-cost scheduling solution that meets all the above constraints while fulfilling customer demand. Associated paramters are detailed in “parameters.csv” Ground-Truth Model (Label). Indices. $i\in I=\{1,\dots,10\}$ (trucks), $t\in T=\{1,\dots,4\}$ (periods). Variables. $w_{it}\geq 0$, $y_{it}\in\{0,1\}$, $u_{it}\in\{0,1\}$ for $t\geq 2$. Objective. $\min\sum_{i\in I}\sum_{t=1}^{4}C_{i}w_{it}\;+\;\sum_{i\in I}\Big(S_{i}y_{i1}+\sum_{t=2}^{4}S_{i}u_{it}\Big).$ Constraints. (1) $\sum_{i\in I}w_{it}\geq d_{t},$ $t\in T,$ (2) $w_{it}\leq Q_{i}y_{it},$ $i\in I,\ t\in T,$ (3) $\sum_{i\in I}w_{it}\leq 0.9\sum_{i\in I}Q_{i}y_{it},$ $t\in T,$ (4) $u_{it}\geq y_{it}-y_{i,t-1},$ $i\in I,\ t=2,3,4,$ $u_{it}\leq y_{it},\qquad u_{it}\leq 1-y_{i,t-1},$ (5) $y_{i1}\leq y_{i2},$ $i\in I,$ $y_{it}+y_{i,t+1}\geq 2u_{it},$ $i\in I,\ t=2,3,$ $u_{i4}=0,$ $i\in I,$ (6) $y_{i,t-1}-y_{it}\leq 1-y_{i,t+1},$ $i\in I,\ t=2,3,$ (7) $|w_{it}-w_{i,t-1}|\leq 300,$ $i\in I,\ t=2,3,4.$ Model Generated by Lean (GPTOSS-20b). Variables. $x_{it}\geq 0$, $y_{it}\in\{0,1\}$, $s_{it}\in\{0,1\}$, $z_{it}\in\{0,1\}$. Objective. $\min\sum_{i\in I}\sum_{t=1}^{4}\big(C_{i}x_{it}+S_{i}s_{it}\big),\qquad s_{i4}=0.$ Constraints. (1) $\sum_{i\in I}x_{it}\geq d_{t},$ $t\in T,$ (2) $x_{it}\leq Q_{i}y_{it},$ $i\in I,\ t\in T,$ (3) $\sum_{i\in I}x_{it}\leq 0.9\sum_{i\in I}Q_{i}y_{it},$ $t\in T,$ (4) $s_{it}\geq y_{it}-y_{i,t-1},\ s_{it}\leq y_{it},$ $s_{it}\leq 1-y_{i,t-1},$ (5) $y_{i,t+1}\geq s_{it},$ $i\in I,\ t=1,2,$ $s_{i4}=0,$ (6) $z_{it}\geq y_{i,t-1}-y_{it},\ z_{it}\leq y_{i,t-1},$ $z_{it}\leq 1-y_{it},\ y_{i,t+1}\leq 1-z_{it},$ $y_{i,t+2}\leq 1-z_{it},$ $i\in I,\ t=1,2,$ (7) $|x_{it}-x_{i,t-1}|\leq 300,$ $i\in I,\ t=2,3,4,$ (+) $\sum_{i\in I}y_{it}=3,$ $t\in T.$ Key Differences. Constraint (+). The generated model imposes an extra requirement that exactly three trucks must be active in each period, $\sum_{i\in I}y_{it}=3$ for all $t\in T$. This constraint is not present in the intended formulation and directly alters the feasible set. Constraint (5). In the intended model, a startup at $t=2,3$ enforces operation over two consecutive periods. In contrast, the coded model only guarantees continuation for startups at $t=1,2$, leading to a weaker minimum-runtime condition. Constraint (6). While the intended formulation enforces a one-period cooldown after shutdown, the generated stop variables may induce a longer forced-off pattern, thereby further restricting feasible operating schedules. Objective and Constraint (4). Startup costs are charged via $s_{it}$ (with $s_{i4}=0$) in the generated model, rather than through $y_{i1}$ and $u_{it}\ (t\geq 2)$ as in the intended formulation. Equivalence holds only if $s_{it}$ perfectly replicates the semantics of the original startup variables. Although the generated formulation differs from the label model in several structural aspects, the two models still yield identical optimal solutions under certain parameter realizations. In particular, when the additional or altered constraints are non-binding at the optimum, the optimal decision vector lies in the intersection of the feasible regions of both formulations.
Appendix I Example Problem Instance in Air-NRM-CA